# Analog Spiking Neuron in CMOS 28 nm Towards Large-Scale Neuromorphic Processors

### v1

@@ VI Conclusion
The research also underscores the potential of CMOS technology in neuromorphic computing by comparing various spiking neuron implementations, highlighting the trend toward more efficient designs. The proof-of-concept SNN, trained on the MNIST dataset, with reduced input data size and quantized weights, emphasizes the feasibility of employing biomimetic models for efficient hardware adaptations, even with low-precision parameters. Overall, the findings from this paper suggest that the proposed CMOS implementation of the LIF spiking neuron holds great promise for advancing neuromorphic computing, offering a viable pathway for creating efficient, scalable, and robust neuromorphic systems.
@@ Acknowledgments
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.
Authors Contributions
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.

### v2

@@ VI Conclusion
The research also underscores the potential of CMOS technology in neuromorphic computing by comparing various spiking neuron implementations, highlighting the trend toward more efficient designs. The proof-of-concept SNN, trained on the MNIST dataset, with reduced input data size and quantized weights, emphasizes the feasibility of employing biomimetic models for efficient hardware adaptations, even with low-precision parameters. Overall, the findings from this paper suggest that the proposed CMOS implementation of the LIF spiking neuron holds great promise for advancing neuromorphic computing, offering a viable pathway for creating efficient, scalable, and robust neuromorphic systems. The current design demonstrates the feasibility of hardware-based SNNs but is limited by the absence of inhibitory mechanisms and the precision of parameter control. Future work will focus on addressing these challenges to improve network performance and further validate the neuron design in more complex tasks.
Future work will focus on utilizing the proposed LIF neuron circuit in networks employing efficient coding schemes, such as spike-latency or event-based coding, to fully leverage the energy-saving and computational advantages of spiking neural networks.
@@ Acknowledgments
The authors would like to acknowledge and thank the Canadian Microsystems Corporation (CMC) and the Natural Sciences and Engineering Research Council (NSERC) for their financial support of this project and for providing the infrastructure essential for ASIC fabrication.